OpenStudy (anonymous):

Factor 54x^4+9x^3+12x^2+2x completely.

OpenStudy (anonymous):

9x^3(6x+1)+2x(6x+1)

OpenStudy (anonymous):

thanks:D

OpenStudy (anonymous):

You're welcome!

Directrix (directrix):

@graceburns9 --> This is not a factorization: 9x^3(6x+1)+2x(6x+1). If I said factor 6 and you said 4 + 2, that would be incorrect because 4 times 2 is not 6. Likewise, 9x^3(6x+1) times 2x(6x+1) is not 54x^4+9x^3+12x^2+2x. There is more work to do here to get the factorization.

OpenStudy (anonymous):

But how do you factor this, I completely forgot how to factor it

Directrix (directrix):

54x^4+9x^3+12x^2+2x = 9x^3(6x+1)+2x(6x+1) = (Factor out common factor of (6x+1) (6x + 1) (9x^3 + 2x ) = (6x + 1) * x * (9 x^2 + 2)= (Factor out common x) x* (6x + 1)* (9 x^2 + 2) @graceburns9

OpenStudy (anonymous):

thanks for takign the pateince to help em out i really appreciate it:)!!